A string is stretched to a length of 260 cm and
both ends are fixed.
If the density of the string is 0.019 g/cm,
and its tension is 263 N, what is the fundamental frequency?
Answer in units of Hz.
1
Expert's answer
2020-02-21T10:05:42-0500
"f=\frac{v}{\lambda}=\frac{1}{2L}\sqrt{\frac{T}{\mu}}"

"f=\frac{1}{2(2.6)}\sqrt{\frac{263}{0.0019}}=71.5\ Hz"